40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 25 22:49:13 2011, Koji, Update, CDS, Front End multiple crash 
    Reply  Wed Jan 26 10:58:48 2011, josephb, Update, CDS, Front End multiple crash 
       Reply  Wed Jan 26 12:03:45 2011, Koji, Update, CDS, Front End multiple crash 
Message ID: 4206     Entry time: Wed Jan 26 10:58:48 2011     In reply to: 4203     Reply to this: 4207
Author: josephb 
Type: Update 
Category: CDS 
Subject: Front End multiple crash 

Looking at dmesg on c1lsc, it looks like the model is starting, but then eventually times out due to a long ADC wait. 

[  114.778001] c1lsc: cycle 45 time 23368; adcWait 14; write1 0; write2 0; longest write2 0
[  114.779001] c1lsc: ADC TIMEOUT 0 1717 53 181

I'm not sure what caused the time out, although there about 20 messages indicating a failed time stamp read from c1sus (its sending TRX information to c1lsc via the dolphin connection) before the time out.

Not seeing any other obvious error messages, I killed the dead c1lsc model by typing:

sudo rmmod c1lscfe

I then tried starting just the front end model again by going to the /opt/rtcds/caltech/c1/target/c1lsc/bin/ directory and typing:

sudo insmod c1lscfe.ko

This started up just the FE again (I didn't use the restart script because the EPICs processes were running fine since we had non-white channels).  At the moment, c1lsc is now running and I see green lights and 0x0 for FB0 status  on the C1LSC_GDS_TP screen.

At this point I'm not sure what caused the timeout.  I'll be adding some more trouble shooting steps to the wiki though.  Also, c1scx, c1scy are probably in need of restart to get them properly sync'd to the framebuilder.

I did a quick test on dataviewer and can see LSC channels such as C1:LSC-TRX_IN1, as well other channels on C1SUS such as BS sensors channels.

Quote:

STATUS:

  • Rebooted c1lsc and c1sus. Restarted fb many times.
  • c1sus seems working.
  • All of the suspensions are damped / Xarm is locked by the green
  • Thermal control for the green is working
  • c1lsc is frozen
  • FB status: c1lsc 0x4000, c1scx/c1scy 0x2bad
  • dataviewer not working 

 

ELOG V3.1.3-